OpenStudy (anonymous):

Write the augmented matrix, and then solve. [3y=15-9x -6x=2y+10

OpenStudy (anonymous):

A. (15, 10) B. (1/4, -1/20) C. no solution D. (3, -2)

OpenStudy (inkyvoyd):

Well first, write it in it's matrix form.

OpenStudy (anonymous):

How would you do that? Can you please explain how to do that?

OpenStudy (anonymous):

CAN SOMEONE PLEASE HELP!!!!!!!!!!

OpenStudy (inkyvoyd):

Write the equation in the form ax+by=c dx+ey=f

OpenStudy (anonymous):

Practice writing augmented matrix first. Simply move terms around so that all the x are in one column all the y are in another columns and all the real numbers are in a different.

OpenStudy (anonymous):

Okay, and after that?

OpenStudy (anonymous):

Assuming you are familiar with the terms and steps. Row reduce the augmented matrix to its reduced echelon form using row operations.

OpenStudy (anonymous):

Did you understand that?

OpenStudy (anonymous):

okay so would my answer be A. (15, 10) ?

OpenStudy (anonymous):

I got something that is not there Ill check again.

OpenStudy (anonymous):

Okay, cause these are the only answers that are possible..

OpenStudy (amistre64):

divide the top by 3 and the bottom by -2 and youll notice something

OpenStudy (anonymous):

@Romero Really? Well maybe there are more than one right answer?

OpenStudy (amistre64):

9x+3y=15 , /3 -6x-2y=10 , /-2 3x+y=5 3x+y=-5

OpenStudy (amistre64):

sarah, if all you want is an answer; connie gave you an answer i spose ... but it definantly aint a "right" answer :)

OpenStudy (anonymous):

@amistre64 actually I asked for someone to explain how to do this so I'm not ONLY looking for an answer.

OpenStudy (anonymous):

from 3x+y=5 3x+y=-5 we get the augemented matrix form 3 1 5 3 1 -5 3 1 5 0 0 -10 i added the -row1 to row2 row2 can then be rewritten as 0+0=-10 0=-10 Can you tell me when 0=-10?

OpenStudy (anonymous):

No..I'm not sure how you got that.

OpenStudy (anonymous):

Thanks for your help though..

OpenStudy (anonymous):

Read your book then. This is the best description I can give sorry.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Your book probably has a lot of examples. If you need some more example just send me a message and I can scan some from my book and send them to you.

OpenStudy (anonymous):

THANKS!